# Scaling and Load-Balancing Equi-Joins

### v2

@@ 4.2 Chunking Lists of Hot Keys
$p=\frac{1+\log_{\min(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}(\max(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}{1+2\times\log_{\min(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}(\max(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}$ (1)
@@
To illustrate using an example, assuming two joined lists each has $10^{5}$ records. A naïve single-executor-per-key Shuffle-Join would output ${10^{5}}^{2}$ pairs of records, i.e., $2\times 10^{10}$ IO, using one executor. Meanwhile, the treeJoinBasic splitter would chunk each list into $47$ sub-lists, and would output $47\times 47\times(2128+2128)$ records, which is $\approx 9.4\times 10^{6}$ IO. These can be distributed on up to $2209$ subsequent executors. Each one of these executors would output $2128\times 2128\times 2\approx 9.1\times 10^{6}$ IO. Notice the only overhead for utilizing up to $2209$ executors to process the join of this key is producing $\approx 9.4\times 10^{6}$ IO by the splitter, and sending this data over the network. Since $\frac{9.4\times 10^{6}}{2\times 10^{10}}<0.05\%$, this overhead is $\approx(1+\lambda)\times 0.05\%$. If the lists instead had $10^{4}$ records, the load can be distributed on up to $484$ executors, and the overhead would only increase to $(1+\lambda)\times 0.2\%$. If the lists instead had $10^{3}$ records, the load can be distributed on up to $100$ executors, and the overhead would only increase to $(1+\lambda)\times 1\%$.
4.3 The Fully-Balanced Tree-Join Algorithm
The treeJoinBasic algorithm (§ 4.1) and optimal list-chunking (§ 4.2) achieve load balancing and good resource utilization. However, this is only true once the initial joined index has been built, and the first-iteration splitters are done splitting the joined lists of the hot keys (Lines $1$ and $4$ of Alg. 1). However, these early operations suffer from load imbalance, since early executors handle complete lists of hot keys. A first-iteration executor that is assigned a hot key may (a) have higher splitting load than other splitters in the first-iteration, and (b) not be able to fit all the records of that hot key in its memory to perform the splitting, resulting in thrashing and hence throttling the join [84]. To achieve maximal load balancing among the executors building the initial joined index, and performing the first-iteration splitting, we identify hot keys, and we distribute the splitting of the initial joined index across a grid of partitions. The distribution of the splitting is similar to the mapRec function in the multi-executor-per-key Shuffle-Join algorithms [2, 65, 15, 56], albeit working at the key level, instead of the executor level.
@@ 4.3 The Fully-Balanced Tree-Join Algorithm
As a preprocessing step, hot keys are identified using the approximate distributed heavy-hitters algorithm in [3]. We only discuss $\mathcal{R}$, but the logic also applies equally to $\mathcal{S}$. The algorithm runs the local Space-Saving algorithms [61, 62] on individual $\mathcal{R}$ partitions, and then merges the local results. We can do the merging over the network in a tree-like manner using a priority queue of bounded size, $|\kappa_{\mathcal{R}}|_{max}$, where $|\kappa_{\mathcal{R}}|_{max}$ is the number of hot keys to be collected from $\mathcal{R}$ [^7] Quantifying $|\kappa_{\mathcal{R}}|_{max}$ and $|\kappa_{\mathcal{S}}|_{max}$ based on multiple parameters pertaining to the datasets and the hardware is discussed in § 7.. If $\mathcal{R}$ is already partitioned by the key, all records of any key reside on one partition, and its local frequency is equal to its global frequency. In that case, exact local counting and merging over the network with a bounded-size priority queue identifies the global hot keys exactly.
@@
10:   ($\ell_{\mathcal{S}},\ell_{\mathcal{R}}$) = $\kappa_{\mathcal{R}\mathcal{S}}$[$key_{rec_{i}}$]
@@
12:   ($\ell_{\mathcal{R}},\ell_{\mathcal{S}}$) = $\kappa_{\mathcal{R}\mathcal{S}}$[$key_{rec_{i}}$]
@@
Notice that load balancing based on augmented-keys assumes the records of the hot keys are randomly distributed among the input partitions of $\mathcal{R}$ and $\mathcal{S}$. If this is not the case, the records of the hot keys can be migrated and balanced among the input partitions. Detecting high variance in the hot key frequencies across partitions can be done while merging their frequencies in a tree-like manner. A similar approach has been employed in [68, 69] to minimize the communication cost in Track-Join, a PRPD-like algorithm.
Figure 5: A schematic example for joining a specific hot key, $k_{1}$, using the load-balanced Tree-Join algorithm. (a) $\mathcal{R}_{H}$ and $\mathcal{S}_{H}$, the respective hot splits of the $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The $k_{1}$ records replicated among multiple augmented $k_{1}$ keys in the unraveled forms of $\mathcal{R}_{H}$ and $\mathcal{S}_{H}$, $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$, respectively. (c) The records of the unravelled relations grouped by the augmented keys using groupByKey to form the first joined index, $joined\_index_{AK}$.
@@
Figure 6: An example showing the difference in load balancing in the first iteration between treeJoinBasic and treeJoin when joining the $\mathcal{R}$ and $\mathcal{S}$ given in Fig. 4. The example assumes each dataset has $6$ partitions. (a) The partitions of the input relations. (b) The partitions of the initial index, $joined\_index_{0}$, built from the input relations in (a) using treeJoinBasic. (c) through (f) show the processing steps of treeJoin. (c) The $\kappa_{\mathcal{R}\mathcal{S}}$ broadcasted to all executors containing the hot-keys and their frequencies in $\mathcal{R}$ and $\mathcal{S}$. (d) The partitions of the cold joined index. (e) The partitions of $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$. (f) The partitions of $joined\_index_{AK}$ built using $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$.
@@
4.4 Executing Natural Self-Joins
@@ 4.5 Defining Hot Keys
$\ell^{2}>\left((1+\lambda)+\left\lceil\frac{\ell^{\frac{2}{3}}}{n}\right\rceil\right)\times\ell^{\frac{4}{3}}$ (3)
@@
4.6 The Number of Iterations
Since the overhead of allocating the executors for a stage was ignored, we make the case the number of iterations in this multistage join is very small.
@@ 4.6 The Number of Iterations
$t<\log_{\frac{3}{2}}\left(\log_{1+\lambda}(\ell_{max})\right)-1$ (4)
@@
4.7 Comparison with the State-of-the-Art Algorithms
The only join algorithms that handle keys that are hot on both sides are the multi-executor-per-key Shuffle-Join algorithms [2, 65, 15, 56]. These algorithms distributes the work of one hot key on multiple executors. While these algorithms achieve excellent load balancing, they put immense pressure on the memory of the executors. This high executor memory requirements makes them less scalable when compared to Tree-Join that requires less executor memory. On the other hand, Tree-Join may take longer to compute the join in the case of very skewed datasets, since it executes potentially multiple iterations.
4.7.1 Algorithmic Comparison
Fig. 2, Fig. 3, and Fig. 5 illustrate the difference between multi-executor-per-key Shuffle-Join algorithms [2, 65, 15, 56], the basic Tree-Join algorithm and the load-balanced Tree-Join algorithm. The three figures focus on the join of a specific key, $k_{1}$, in the input datasets.
@@ 4.7.1 Algorithmic Comparison
The multi-executor-per-key Shuffle-Join algorithms produces the join results in one MapReduce job, while the Tree-Join algorithms take multiple iterations. These iterations may take longer time to finish but consumes less memory at each iteration, as shown next in § 4.7.2. Overall this enhances the scalability of Tree-Join.
4.7.2 Memory Requirements Analysis
For simplicity, we assume the input records are distributed randomly on the input dataset partitions. and we ignore the $\lceil$.$\rceil$ . ⌉ $\lfloor$.$\rfloor$ . ⌋ functions. For the sake of discussion, we focus on $\mathcal{R}$-$\mathcal{S}$ equi-joins, but the discussion also applies equally to natural self-joins [65, 56]. We also assume in-memory processing, where the input and output of each MapReduce function have to fit in memory during processing. That is, for equi-joins, the records input to and output by any executor for any given key have to fit in the memory of that executor.
We compare with the most recent algorithm from the multi-executor-per-key Shuffle-Join family of algorithms. ExpVar-Join [56] tries to find the optimal grid dimensions for each hot key that would balance the load on all executors.
@@ 4.7.2 Memory Requirements Analysis
5 The Small-Large Outer-Joins
@@
The scenario of Small-Large joins arise when $\mathcal{R}\gg\mathcal{S}$ and $\mathcal{S}$ can be assumed to fit in the memory of each executor. An index can be built for the small relation, $\mathcal{S}$, the driver collects the index locally using the treeAggregate operation that is implemented as a series of MapReduce jobs. The driver then broadcasts the index to all $n$ executors performing the join. Broadcasting $\mathcal{S}$ can be done in time that is logarithmic in $n$ [12, 75] [^9] The tree-based broadcast algorithms streams the data from the driver to a layer of executors. The data is then streamed from these executors to a wider layer of executors, and so on. So, each executor acts as a receiver and a sender. Multiple concurrent such trees could be used to optimally utilizes the entire bandwidth of the network, and hence finish the broadcast in $O(n)$ rounds of communication., and can hence be faster than shuffling the large relation, $\mathcal{R}$, across the network.
Figure 7: Synthesizing the results of the full-outer-join from other joins. (a) shows the input relations whose full-outer-join results, shown in (b), can be expressed as a union of the left-outer-join results in (c) and the right-anti-join results in (d). The shaded areas represent keys whose records are included in the join results.
@@ 5.1 Optimizations
  4. 4.
    In a shared-nothing architecture that supports both multicasting and broadcasting, instead of broadcasting the index to all the executors (Line $2$ in Alg. 13 and Alg. 16), the driver may broadcast the unique keys only. Instead of the local joins (Line $3$ in Alg. 13 and Alg. 16), each executor, $e_{i}$, joins the $\mathcal{S}$ keys with its $\mathcal{R}_{i}$ and sends the joinable keys to the driver. The driver multicasts each $\mathcal{S}$ record only to the executors that have its joinable partitions. Each $e_{i}$ joins the received records with its $\mathcal{R}_{i}$. The driver need not compute the unjoinable $\mathcal{S}$ keys (Lines 44448888 in Alg. 16), since they are the keys not multicasted at all. This optimization is a form of semi-join reduction, which was first introduced in [16, 7], and later revived in other optimizations, e.g., [81, 47].
5.2 Comparison with the State-of-the-Art Algorithms
A main difference between the proposed IB-FO-Join (Alg. 16), DER [91] and DDR [27] is subtle but effective. To flag the unjoined records in the $n$ executors, IB-FO-Join utilizes the unique keys in a semi-join fashion, instead of the DBMS-assigned record ids in case of DER [^10] A minor advantage of the IB-Join family is abolishing the dependency on DBMS-specific functionality. This facilitates its adoption in any shared-nothing architecture. or entire records in case of DDR. Assuming the key is of the same size as the record id, and is significantly smaller than the record, sending unique keys reduces the network load, and scales better with a skewed $\mathcal{S}$.
@@ 6 The AM-Join Algorithm
Figure 8: An example for joining the two relations in Fig. 1 using AM-Join. (a) The hot-key and cold-key sub-relations of $\mathcal{R}$ and $\mathcal{S}$ after the first split. (b) The four sub-relations for each of the two input relations after the second split, and their joins based on Eqn. 5.
@@
Since the first join involves keys that are hot in both the joined relations, Tree-Join (Alg. 10) is used. The second and the third joins in Eqn. 5 lend themselves naturally to Small-Large joins. We only discuss the second join, but the logic also applies equally to the third join. $\mathcal{R}_{HC}$ and $\mathcal{S}_{CH}$ contain the $\mathcal{R}$ and $\mathcal{S}$ records, respectively, with keys that are hot in $\mathcal{R}$ but are cold in $\mathcal{S}$. Hot keys are rare by definition, since a hot key requires a minimum number of records to be hot, i.e., a minimum percentage of the records in the relation. Hence, the number of keys in $\mathcal{R}_{HC}$ and $\mathcal{S}_{CH}$ is limited. Moreover, the number of records with these keys in $\mathcal{S}$ is limited, since these keys are cold in $\mathcal{S}$. Therefore, $\mathcal{S}_{CH}$ is bounded in size, and should fit in the memory of the executors. This is discussed in more details in § 7.1, where tight bounds on the number of hot keys are established to accommodate the smaller relation in the memory of the executors.
@@
6.1 Splitting the Relations
Once the hot keys for both $\mathcal{R}$ and $\mathcal{S}$ are collected, each relation is split into its four sub-relations. We only discuss $\mathcal{R}$, but the logic also applies equally to $\mathcal{S}$. This splitRelation module (Alg. 22) proceeds in two rounds. It first splits $\mathcal{R}$ into $\mathcal{R}_{H}$, the sub-relation whose keys are in $\kappa_{\mathcal{R}}$, and $\mathcal{R}_{C}$, the cold-key records in $\mathcal{R}$. In the example in Fig. 1, the results of the first-round of splitting is in Fig. 8(a). In the second round, $\mathcal{R}_{H}$ is split into $\mathcal{R}_{HH}$, the sub-relation whose keys are in $\kappa_{\mathcal{S}}$, and $\mathcal{R}_{HC}$, the remaining records in $\mathcal{R}_{H}$. $\mathcal{R}_{C}$ is similarly split using $\kappa_{\mathcal{S}}$ into $\mathcal{R}_{CH}$ and $\mathcal{R}_{CC}$. In the example in Fig. 1, the second-round of splitting are in Fig. 8(b).

### v3

@@ 4.2. Chunking Lists of Hot Keys
(1) $p=\frac{1+\log_{\min(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}(\max(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}{1+2\times\log_{\min(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}(\max(\ell_{\mathcal{R}},\ell_{\mathcal{S}}))}$
@@
To illustrate using an example, assuming two joined lists, each has $10^{5}$ records. A naïve single-executor-per-key Shuffle-Join would output ${10^{5}}^{2}$ pairs of records, i.e., $2\times 10^{10}$ IO, using one executor. Meanwhile, the treeJoinBasic splitter would chunk each list into $47$ sub-lists, and would output $47\times 47\times(2128+2128)$ records, which is $\approx 9.4\times 10^{6}$ IO. These can be distributed on up to $2209$ subsequent executors. Each one of these executors would output $2128\times 2128\times 2\approx 9.1\times 10^{6}$ IO. Notice the only overhead for utilizing up to $2209$ executors to process the join of this key is producing $\approx 9.4\times 10^{6}$ IO by the splitter, and sending this data over the network. Since $\frac{9.4\times 10^{6}}{2\times 10^{10}}<0.05\%$, this overhead is $\approx(1+\lambda)\times 0.05\%$. If the lists instead had $10^{4}$ records, the load can be distributed on up to $484$ executors, and the overhead would only increase to $(1+\lambda)\times 0.2\%$. If the lists instead had $10^{3}$ records, the load can be distributed on up to $100$ executors, and the overhead would only increase to $(1+\lambda)\times 1\%$.
4.3. The Fully-Balanced Tree-Join Algorithm
The treeJoinBasic algorithm (§ 4.1) and optimal list-chunking (§ 4.2) achieve load balancing and good resource utilization. However, this is only true once the initial joined index has been built, and the first-iteration splitters are done splitting the joined lists of the hot keys (Lines $1$ and $4$ of Alg. 1). However, these early operations suffer from load imbalance, since early executors handle complete lists of hot keys. A first-iteration executor that is assigned a hot key may (a) have higher splitting load than other splitters in the first-iteration, and (b) not be able to fit all the records of that hot key in its memory to perform the splitting, resulting in thrashing and hence throttling the join (Suri and Vassilvitskii, 2011). To achieve maximal load balancing among the executors building the initial joined index, and performing the first-iteration splitting, we identify hot keys, and we distribute the splitting of the initial joined index across a grid of partitions. The distribution of the splitting is similar to the mapRec function in the multi-executor-per-key Shuffle-Join algorithms (Afrati and Ullman, 2010; Okcan and Riedewald, 2011; Beame et al., 2014; Li et al., 2018), albeit working at the key level, instead of the executor level.
As a preprocessing step, hot keys are identified using the approximate distributed heavy-hitters algorithm in (Agarwal et al., 2013). We only discuss $\mathcal{R}$, but the logic also applies equally to $\mathcal{S}$. The algorithm runs the local Space-Saving algorithms (Metwally et al., 2005, 2006) on individual $\mathcal{R}$ partitions, and then merges the local results. We can do the merging over the network in a tree-like manner using a priority queue of bounded size, $|\kappa_{\mathcal{R}}|_{max}$, where $|\kappa_{\mathcal{R}}|_{max}$ is the number of hot keys to be collected from $\mathcal{R}$ [^7] Quantifying $|\kappa_{\mathcal{R}}|_{max}$ and $|\kappa_{\mathcal{S}}|_{max}$ based on multiple parameters pertaining to the datasets and the hardware is discussed in § 7.. If $\mathcal{R}$ is already partitioned by the key, all records of any key reside on one partition, and its local frequency is equal to its global frequency. In that case, exact local counting and merging over the network with a bounded-size priority queue identifies the global hot keys exactly.
@@ 4.3. The Fully-Balanced Tree-Join Algorithm
10:   $\langle\ell_{\mathcal{S}},\ell_{\mathcal{R}}\rangle$ = $\kappa_{\mathcal{R}\mathcal{S}}$[$key_{rec_{i}}$]
@@
12:   $\langle\ell_{\mathcal{R}},\ell_{\mathcal{S}}\rangle$ = $\kappa_{\mathcal{R}\mathcal{S}}$[$key_{rec_{i}}$]
@@
Notice that load balancing based on augmented-keys assumes the records of the hot keys are randomly distributed among the input partitions of $\mathcal{R}$ and $\mathcal{S}$. If this is not the case, the records of the hot keys can be migrated and balanced among the input partitions. Detecting high variance in the hot key frequencies across partitions can be done while merging their frequencies in a tree-like manner. A similar approach has been employed in (Polychroniou et al., 2014, 2018) to minimize the communication cost in Track-Join, a PRPD-like algorithm.
Figure 5. A schematic example for joining a specific hot key, $k_{1}$, using the load-balanced Tree-Join algorithm. (a) $\mathcal{R}_{H}$ and $\mathcal{S}_{H}$, the respective hot splits of the $\mathcal{R}$ and $\mathcal{S}$ relations to be joined. (b) The $k_{1}$ records replicated among multiple augmented $k_{1}$ keys in the unraveled forms of $\mathcal{R}_{H}$ and $\mathcal{S}_{H}$, $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$, respectively. (c) The records of the unravelled relations grouped by the augmented keys using groupByKey to form the first joined index, $joined\_index_{AK}$.
@@
Figure 6. An example showing the difference in load balancing in the first iteration between treeJoinBasic and treeJoin when joining the $\mathcal{R}$ and $\mathcal{S}$ given in Fig. 4. The example assumes each dataset has $6$ partitions. (a) The partitions of the input relations. (b) The partitions of the initial index, $joined\_index_{0}$, built from the input relations in (a) using treeJoinBasic. (c) through (f) show the processing steps of treeJoin. (c) The $\kappa_{\mathcal{R}\mathcal{S}}$ broadcasted to all executors containing the hot-keys and their frequencies in $\mathcal{R}$ and $\mathcal{S}$. (d) The partitions of the cold joined index. (e) The partitions of $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$. (f) The partitions of $joined\_index_{AK}$ built using $unraveled_{\mathcal{R}_{H}}$ and $unraveled_{\mathcal{S}_{H}}$.
@@
4.4. Executing Natural Self-Joins
@@ 4.5. Defining Hot Keys
(3) $\ell^{2}>\left((1+\lambda)+\left\lceil\frac{\ell^{\frac{2}{3}}}{n}\right\rceil\right)\times\ell^{\frac{4}{3}}$
@@
4.6. The Number of Iterations
Since the overhead of allocating the executors for a stage was ignored, we make the case that the number of iterations in this multistage join is very small.
@@ 4.6. The Number of Iterations
(4) $t<\log_{\frac{3}{2}}\left(\log_{1+\lambda}(\ell_{max})\right)-1$
@@
4.7. Comparison with the State-of-the-Art Algorithms
The only join algorithms that handle keys that are hot on both sides are the multi-executor-per-key Shuffle-Join algorithms (Afrati and Ullman, 2010; Okcan and Riedewald, 2011; Beame et al., 2014; Li et al., 2018). These algorithms distribute the work of one hot key on multiple executors. While these algorithms achieve excellent load balancing, they put immense pressure on the memory of the executors. This high executor memory requirements makes them less scalable when compared to Tree-Join that requires less executor memory. On the other hand, Tree-Join may take longer to compute the join in the case of very skewed datasets, since it executes potentially multiple iterations.
4.7.1. Algorithmic Comparison
Fig. 2, Fig. 3, and Fig. 5 illustrate the difference between multi-executor-per-key Shuffle-Join algorithms (Afrati and Ullman, 2010; Okcan and Riedewald, 2011; Beame et al., 2014; Li et al., 2018), the basic Tree-Join algorithm and the load-balanced Tree-Join algorithm. The three figures focus on the join of a specific key, $k_{1}$, in the input datasets.
@@ 4.7.1. Algorithmic Comparison
The multi-executor-per-key Shuffle-Join algorithms produce the join results in one MapReduce job, while the Tree-Join algorithms take multiple iterations. These iterations may take longer time to finish but consume less memory at each iteration, as shown next in § 4.7.2. Overall this enhances the scalability of Tree-Join.
4.7.2. Memory Requirements Analysis
For simplicity, we assume the input records are distributed randomly on the input dataset partitions. and we ignore the $\lceil$.$\rceil$ . ⌉ $\lfloor$.$\rfloor$ . ⌋ functions. For the sake of discussion, we focus on $\mathcal{R}$-$\mathcal{S}$ equi-joins, but the discussion also applies equally to natural self-joins (Okcan and Riedewald, 2011; Li et al., 2018). We also assume in-memory processing, where the input and output of each MapReduce function have to fit in memory during processing. That is, for equi-joins, the records input to and output by any executor for any given key have to fit in the memory of that executor.
We compare the Tree-Join algorithms with the most recent algorithm from the multi-executor-per-key Shuffle-Join family of algorithms. ExpVar-Join (Li et al., 2018) tries to find the optimal grid dimensions for each hot key that would balance the load on all executors.
@@ 4.7.2. Memory Requirements Analysis
5. The Small-Large Outer-Joins
@@
The scenario of Small-Large joins arise when $\mathcal{R}\gg\mathcal{S}$ and $\mathcal{S}$ can be assumed to fit in the memory of each executor. An index can be built for the small relation, $\mathcal{S}$, the driver collects the index locally using the treeAggregate operation that is implemented as a series of MapReduce jobs. The driver then broadcasts the index to all $n$ executors performing the join. Broadcasting $\mathcal{S}$ can be done in time that is logarithmic in $n$ (Bar-Noy and Kipnis, 1994; Sanders et al., 2009) [^9] The tree-based broadcast algorithms stream the data from the driver to a layer of executors. The data is then streamed from these executors to a wider layer of executors, and so on. So, each executor acts as a receiver and a sender. Multiple concurrent such trees could be used to optimally utilize the entire bandwidth of the network, and hence finish the broadcast in $O(n)$ rounds of communication., and can hence be faster than shuffling the large relation, $\mathcal{R}$, across the network.
Figure 7. Synthesizing the results of the full-outer-join from other joins. (a) shows the input relations whose full-outer-join results, shown in (b), can be expressed as a union of the left-outer-join results in (c) and the right-anti-join results in (d). The shaded areas represent keys whose records are included in the join results.
@@ 5.1. Optimizations
  4. (4)
    In a shared-nothing architecture that supports both multicasting and broadcasting, instead of broadcasting the index to all the executors (Line $2$ in Alg. 13 and Alg. 16), the driver may broadcast the unique keys only. Instead of the local joins (Line $3$ in Alg. 13 and Alg. 16), each executor, $e_{i}$, joins the $\mathcal{S}$ keys with its $\mathcal{R}_{i}$ and sends the joinable keys to the driver. The driver multicasts each $\mathcal{S}$ record only to the executors that have its joinable partitions. Each $e_{i}$ joins the received records with its $\mathcal{R}_{i}$. The driver need not compute the unjoinable $\mathcal{S}$ keys (Lines 44448888 in Alg. 16), since they are the keys not multicasted at all. This optimization is a form of semi-join reduction, which was first introduced in (Bernstein et al., 1981; Apers et al., 1983), and later revived in other optimizations, e.g., (Stocker et al., 2001; Hill and Ross, 2009).
5.2. Comparison with the State-of-the-Art Algorithms
A main difference between the proposed IB-FO-Join (Alg. 16), DER (Xu and Kostamaa, 2010) and DDR (Cheng et al., 2017) is subtle but effective. To flag the unjoined records in the $n$ executors, IB-FO-Join utilizes the unique keys in a semi-join fashion, instead of the DBMS-assigned record ids in case of DER [^10] A minor advantage of the IB-Join family is abolishing the dependency on DBMS-specific functionality. This facilitates its adoption in any shared-nothing architecture. or entire records in case of DDR. Assuming the key is of the same size as the record id, and is significantly smaller than the record, sending unique keys reduces the network load, and scales better with a skewed $\mathcal{S}$.
@@ 6. The AM-Join Algorithm
Figure 8. An example for joining the two relations in Fig. 1 using AM-Join. (a) The hot-key and cold-key sub-relations of $\mathcal{R}$ and $\mathcal{S}$ after the first split. (b) The four sub-relations for each of the two input relations after the second split, and their joins based on Eqn. 5.
@@
Since the first join involves keys that are hot in both the joined relations, Tree-Join (Alg. 10) is used. The second and the third joins in Eqn. 5 lend themselves naturally to Small-Large joins. We only discuss the second join, but the logic also applies equally to the third join. $\mathcal{R}_{HC}$ and $\mathcal{S}_{CH}$ contain the $\mathcal{R}$ and $\mathcal{S}$ records, respectively, with keys that are hot in $\mathcal{R}$ but are cold in $\mathcal{S}$. Hot keys are rare by definition, since a hot key requires a minimum number of records to be hot, i.e., a minimum percentage of the records in the relation. Hence, the number of keys in $\mathcal{R}_{HC}$ and $\mathcal{S}_{CH}$ is limited. Moreover, the number of records with these keys in $\mathcal{S}$ is limited, since these keys are cold in $\mathcal{S}$. Therefore, $\mathcal{S}_{CH}$ is bounded in size, and should fit in the memory of the executors. This is discussed in more detail in § 7.1, where tight bounds on the number of hot keys are established to accommodate the smaller relation in the memory of the executors.
@@
6.1. Splitting the Relations
Once the hot keys for both $\mathcal{R}$ and $\mathcal{S}$ are collected, each relation is split into its four sub-relations. We only discuss $\mathcal{R}$, but the logic also applies equally to $\mathcal{S}$. This splitRelation module (Alg. 22) proceeds in two rounds. It first splits $\mathcal{R}$ into $\mathcal{R}_{H}$, the sub-relation whose keys are in $\kappa_{\mathcal{R}}$, and $\mathcal{R}_{C}$, the cold-key records in $\mathcal{R}$. In the example in Fig. 1, the results of the first-round of splitting is in Fig. 8(a). In the second round, $\mathcal{R}_{H}$ is split into $\mathcal{R}_{HH}$, the sub-relation whose keys are in $\kappa_{\mathcal{S}}$, and $\mathcal{R}_{HC}$, the remaining records in $\mathcal{R}_{H}$. $\mathcal{R}_{C}$ is similarly split using $\kappa_{\mathcal{S}}$ into $\mathcal{R}_{CH}$ and $\mathcal{R}_{CC}$. In the example in Fig. 1, the second-round of splitting is in Fig. 8(b).